# “When to Hand Off, When to Work Together”: Expanding Human-Agent Co-Creative Collaboration through Concurrent Interaction

### 06 Mar 2026 v2

@@ 5.1. Design Implications for Collaborative Agents in Co-Creative Work
Enable temporal coordination through execution milestones that reveal synchronization points for intervention, delegation, and parallel work.
Process transparency’s value lies not just in visibility but also in establishing temporal landmarks to synchronize user actions with agent progress. Observational monitoring appeared in 68.7% of all turns, with 14.0% consisting solely of observation when users lacked mental models. Making execution visible through step-by-step operations enabled users to reason about the agent’s approach, anticipate next actions, and identify optimal intervention timing—P7 and P8 monitored subtask completions to coordinate their parallel work, using operations as synchronization points.
Process transparency directly enabled productive triggers like “Idea Spark from Agent’s Work-in-Progress” (Figure 8), where users observed intermediate outputs and determined when to engage. Participants (P1, P7, P8) reported substantially reduced clarification costs, identifying and correcting misunderstandings mid-execution rather than waiting for complete but incorrect results. While execution speed or mid-output representations(e.g., the level of abstraction (Suh et al., 2024)) can be controlled once users develop confidence, these temporal milestones remain essential for enabling synchronization timing throughout collaboration.
Support multi-level intervention granularity and interpret each level as distinct collaboration signals
@@
Concurrent actions occurred in 31.8% of turns across five intervention levels (Figure 1) signaling different collaboration needs. Fine-grained interventions (demonstration, in-situ co-editing) signaled immediate course correction—P5’s spacing adjustments and P1’s shadow refinements meant “apply this now and going forward.” Mid-level interventions (artifact takeover) indicated parallel work—P2 and P9 duplicated work-in-progress while the agent continued independently. Coarse-grained interventions (appropriation, opportunistic takeover) validated direction—P4 and P7 appropriating outputs meant “this works, continue autonomously.”
The current system treats all concurrent interactions as the same signal. Instead, agents should interpret granularity as calibrated feedback: fine-grained edits demand immediate learning (apply patterns going forward); mid-level takeovers signal task redistribution without behavior change; coarse-grained appropriation confirms direction, warranting increased autonomy. This enables distinguishing “plan correction” from “validation” from “task handoff.” Future systems should propagate learned patterns: apply fine-grained corrections to similar elements within the task; accelerate similar operations after coarse-grained validatio.
Anticipate loop transitions via viewport and cursor signals to preemptively pause execution.
@@
Rather than only reacting after users intervene, agents should track viewport focus and cursor proximity to predict when interventions are imminent and pause proactively. Participants (P2, P4, P10) identified distinct behavioral signals of users that would be helpful to understand the user: when users center their viewport on the agent’s work area (focused observation) while moving their cursor toward the agent’s output, this typically signals (C) Concurrent intervention within seconds. Conversely, when users maintain peripheral viewport (intermittent glances) while keeping their cursor in distant workspace areas, this suggests they will continue (H) Hands-off without intervening.
When detecting imminent intervention signals, agents could preemptively pause execution at the current operation boundary and surface intermediate state for inspection, rather than completing the full planned sequence. For example, if the agent plans to create five color variations but detects intervention signals (focused viewport, approaching cursor) after completing the second variation, it should pause before generating the third—allowing the user to redirect without wasting computation on unwanted variations. Conversely, when detecting hands-off signals (peripheral viewport, distant cursor), agents can proceed with planned execution without artificial pauses.
Sustained observation without loop transition also signals different needs by collaboration stage. Early in collaboration (first 2-3 tasks), sustained observation indicates mental model building—agents should proceed patiently while ensuring operations remain visible. Later, sustained observation without hands-off transition suggests imminent intervention—agents should prepare an intermediate state for inspection and pause at the next operation boundary. This temporal prediction transforms behavioral signals into real-time execution control, enabling agents to synchronize their workflow with anticipated user engagement.
Support direct manipulation of agent presence within the canvas as explicit high-confidence signals.
@@
Participants (P1, P7-P9) expressed a strong desire for such controls, suggesting they would move the agent’s cursor away from their current work area to signal “work independently elsewhere” (Hands-off Loop), or position it near specific elements to indicate “focus your attention here when you do this task” (guiding toward Concurrent or Directive engagement). These spatial manipulations can unambiguously communicate user intent compared to inferring intent from the user’s viewport or observation patterns alone. Direct manipulation enables users to explicitly communicate when (now vs. later), where (which workspace area), and how (proximity suggesting collaboration vs. distance suggesting independence) they expect the agent to work, providing stronger signals for the agent to align its workflow with user expectations. By treating agent presence as a manipulable entity, systems can capture explicit loop transition intent that complements implicit behavioral inference, reducing ambiguity about whether users want autonomous execution, close collaboration, or temporary disengagement.
Synthesize collaboration signals for context-aware proactive behavior.
@@
Current agents operate exclusively in response to user requests, and while recent proactive agents exist, they typically rely on simple heuristics—such as detecting prolonged user inactivity to offer help (Pu et al., 2025)—rather than relying on rich collaboration context. Our findings reveal that effective proactive behavior requires synthesizing diverse signals: loop transitions indicating confidence or concern, spatial manipulations communicating desired autonomy levels, observation patterns revealing engagement, and concurrent interventions showing valued directions.
By integrating these cues, agents can anticipate user needs beyond current requests: recognizing when sustained observation signals uncertainty warranting proactive explanation of current approach; detecting when users consistently delegate certain task types, enabling the agent to proactively suggest similar tasks (“I notice you often delegate color exploration—should I generate variations for this new design?”); or inferring from spatial distancing to work independently on lower-priority subtasks without awaiting explicit instruction. Therefore, future collaborative agents should proactively decide what to work on, how to approach it, and where to focus—not based on rigid heuristics, but by continuously interpreting the collaborative situation through accumulated implicit and explicit signals, enabling truly adaptive co-creative partnership.
5.2. Designing Agent Execution Representations
While the preceding design implications highlight why observable and intervenable execution is critical for collaboration, here we focus on how agent execution should be represented to support such interaction. In our work, making the agent’s execution process visible, interpretable, and interactable was foundational to enabling effective collaboration, including concurrent interaction. However, most current commercial agent systems—except web agents like GPT agent [^8] https://chatgpt.com/features/agent/—only share their text-based reasoning processes rather than showing their execution process. While recent agent systems (e.g., Manus (AI, )) perform diverse tasks quickly and provide key reasoning steps, their text-based explanations pose challenges. Users struggle to understand workflows from text alone, particularly in visually-oriented domains where anticipating outcomes requires seeing intermediate artifacts. Domain-appropriate execution representations are therefore critical: they enable users to quickly grasp “how the agent is working” at a glance, preventing oversight issues and premature “let it finish and see what happens” delegation without process understanding.
Our studies demonstrated that showing execution through familiar design operations enabled participants to identify next tasks, refine directions, catch missing details, and determine parallel work opportunities. Only after users develop understanding through observable, editable processes can true full delegation emerge—not as blind trust, but as informed confidence grounded in the agent’s demonstrated ability to adapt to the user’s working style. Future research should explore domain-specific representation strategies that balance observability, intervention affordance, and execution efficiency.
5.3. Concurrent Interactions as Behavioral Signals for Preference Learning
Among the five concurrent interaction patterns (Figure 1), demonstration-based steering and in-situ co-editing provide rich behavioral signals about user preferences—they reveal how users want work done through direct action rather than verbal specification. Rather than building semantic models of user intent, Cleo treats concurrent interventions as observable behavioral data that signals design directions and quality expectations. When users modify the agent’s work-in-progress or demonstrate alternative approaches, these actions communicate preferences through concrete examples that would be difficult to articulate verbally.
Users’ concurrent modifications on the agent’s work and chosen directions serve as positive signals—“work this way”—while rejected or corrected approaches signal negative feedback—“avoid this direction.” For instance, when P7 took over to apply fine-grained aesthetic refinements, the specific adjustments demonstrated quality expectations and stylistic preferences more precisely than verbal instruction could convey. Similarly, when P1 and P9 appropriated certain design elements while ignoring others, these choices implicitly ranked which directions were worth pursuing.
@@ 5.3. Concurrent Interactions as Behavioral Signals for Preference Learning
Beyond individual interactions, feedback signals emerge from loop transitions themselves (Figure 11). Users moving from observation to concurrent engagement signal specific aspects requiring attention, while transitions to hands-off validate the agent’s approach. Recent research explores generating or simulating user interactions with generative models to produce intent-aligned outputs (Kim et al., 2026). Our characterized interaction patterns, triggers, and loop dynamics could enable more accurate user simulation—providing rich behavioral data about not just what users want, but when, how, and why they intervene, supporting more nuanced intent alignment in future agentic systems.
However, these benefits come with cognitive costs. Monitoring the agent while managing parallel work creates attention burden. Our findings suggest this trade-off varies by task familiarity: when users lacked mental models of the agent’s capabilities (14% observation-only turns), they invested heavily in monitoring; once familiar, users strategically alternated between focused work and periodic checking. Future work should investigate when observation benefits outweigh attention costs and how to minimize monitoring burden through selective notifications or adaptive visibility.
5.4. Extending Concurrent Interaction Beyond Creative Domains
While our findings emerged from design work, the principles of observable execution and concurrent interaction could be extended to other co-creative domains—though each requires domain-specific adaptations.
In text-based domains like coding and writing, agents generate text far faster than humans can comprehend, necessitating multi-scale representation strategies. This allows users to shift between different levels of abstraction. For example, zooming in might reveal full generated text by the agent, while zooming out shows hierarchical summaries—enabling users to grasp direction without reading every word. Visualizing the agent’s current focus (which section, which function) helps users determine intervention timing. Users could provide concurrent feedback through lightweight annotations—marking passages with “keep,” “revise,” or “delete”—capturing implicit stylistic and content preferences without disrupting generation flow.
In web navigation and OS-level tasks, concurrent interaction takes a different form than in artifact-centric creative work. Rather than jointly manipulating the same interface elements, concurrent interaction in these domains is often better supported through resource-level exchange within the agent’s workflow.Users might appropriate currently open webpages from the agent’s workspace, extract specific moments from search history as snapshots, or inject relevant pages into the agent’s path—signaling “consider this” through direct manipulation rather than verbal instruction.
The key to extending concurrent interaction across domains lies in identifying domain-appropriate manipulable objects that enable resource exchange. In design, these are visual elements (frames, layers, components); in coding, they could be code blocks, function definitions, or test cases that users appropriate from agent-generated code; in writing, paragraphs or sections that users extract and reposition; in web navigation, browser tabs, saved pages, or search queries that users inject into the agent’s exploration path. **Concurrent interaction requires reification—making the agent’s intermediate work products into discrete, manipulable objects** rather than ephemeral process states. Future work should investigate which domain-specific artifacts are most valuable as exchange currencies, how to represent them for easy appropriation, and how agents should interpret such exchanges as collaboration signals rather than disruptions.
$5.$ Unique Affordances of Human-Agent Collaboration
@@ 5.4. Extending Concurrent Interaction Beyond Creative Domains
While we draw on human-human collaboration theories  (Gutwin and Greenberg, 2002; Robertson, 2002; Tang, 1991), human-agent collaboration exhibits unique interaction affordances absent in human teams. Notably, users freely terminated agent work (9.7% of turns) or overwrote agent outputs through concurrent takeover—behaviors that would be socially costly or disruptive in human collaboration. This low social cost of correction enables more experimental, iterative workflows: users can interrupt without guilt, redirect without negotiation, and appropriate outputs without attribution concerns.
Participants exploited this asymmetry in ways unthinkable with human collaborators. Participants (P1-P3, P6-P8, and P10) terminated execution when correction felt burdensome, treating the agent as a disposable experiment rather than a collaborator requiring explanation. P5 and P9 appropriated agent outputs without acknowledgment, using agent work as raw material for their own designs.
This low-cost correction dynamic fundamentally differs from human collaboration, where interruption, rejection, or appropriation carry social consequences—potentially damaging relationships, morale, or future collaboration. Agents, by contrast, accept abrupt termination, interpret overwriting as feedback, and restart without loss of motivation. Future work could investigate how this unique dynamic shapes collaboration strategies: do users develop different creative processes with agents versus humans? Does low-cost correction enable more ambitious experimentation or lead to over-reliance on trial-and-error? Understanding these dynamics could inform both agent design and our theoretical models of human-AI collaboration.
@@ 5.6. Limitations and Future Work
While our in-depth analysis of 214 turns provided rich insights, our second study was limited to 10 participants over two days. More in-depth insights into collaboration patterns would require larger-scale, longitudinal validation. Extended usage studies could reveal how interaction patterns evolve as users develop mental models—which loops become dominant or diminish over time, whether delegation increases with familiarity, and how natural usage patterns emerge beyond initial exploration. Such studies would strengthen our decision model’s generalizability and uncover temporal dynamics we could not observe in our short-term study.
Our ReAct-based agent architecture—reasoning before tool calls, self-determining task completion—represents one design choice among many possible agent structures. Different architectures (e.g., reflection-based, hierarchical planning, continuous execution) may elicit different interaction patterns. Investigating how agent architecture influences human collaboration could reveal which design elements are critical for effective co-creative work and inform future agent system design.
Finally, agent execution speed warrants systematic investigation. Slower execution enables users to monitor and comprehend workflows, but faster execution helps users reach complete outcomes earlier. Future work should explore user-controllable speed: when do users want rapid autonomous execution versus step-by-step observable progress? Understanding this preference could inform adaptive systems that adjust pace based on task stage, user confidence, and intervention frequency, balancing efficiency with observability as collaboration matures.
@@ 6. Conclusion
This work investigates the opportunities and considerations of transitioning human-agent collaboration from sequential turn-taking to more dynamic, concurrent interaction within co-creative design. An initial study with the first design probe (N=10) revealed that process transparency helped users understand the agent’s actions and parallelize work, but also demonstrated the need for agents that adapt to users’ concurrent interventions. Thus, we proposed Cleo as the second design probe, a collaborative design agent that interprets and adapts to user actions that occur as it executes its own task. Through a two-day study (N=10), we identified and analyzed 214 turn-level interactions between participants and Cleo—revealing how participants transitioned fluidly from delegation to concurrent co-creation. Furthermore, the study demonstrated that participants applied different patterns of concurrent interaction (e.g., appropriation, takeover, co-editing) based on different situational and collaboration-related factors (e.g., perceived task importance). Finally, this work proposes a set of design implications for designing more responsive and adaptive co-creative agents that support diverse collaboration styles and patterns.

### 07 Apr 2026 v4

@@ 5.1. Design Implications for Collaborative Agents in Co-Creative Work
DI1. Adapt process visibility granularity to user needs and task stage for personalized synchronization. Agents should structure execution as observable milestones that users can use as synchronization points for timing interventions, delegation, and parallel work. Observational monitoring appeared in 68.7% of turns, with P7 and P8 actively timing their own work around the agent’s subtask completions, directly enabling triggers like “Idea Spark from Agent’s Work-in-Progress” (Table 2) and reducing clarification costs by catching misunderstandings mid-execution (P1, P4, and P8). However, rather than exposing progress uniformly, agents should adapt granularity to user needs and task stage: finer-grained operations for high-stakes tasks requiring close oversight (P2, P5, and P9), and larger subplan completions for efficiency-focused stages (P6-P8, and P10). This adaptive granularity (Suh et al., 2024) lets users determine their own synchronization rhythm rather than being locked into a single representation pace.
@@
DI2.Interpret concurrent intervention scope as calibrated signals to tune agent behavior to user workstyle. Concurrent actions occurred in 31.8% of turns across five intervention levels (Table 1), each signaling different collaboration needs. Fine-grained interventions (demonstration, in-situ co-editing) signal immediate course correction, where P5’s spacing adjustments and P1’s shadow refinements meant “apply this now and going forward.” Mid-level interventions (artifact takeover) indicate parallel work, where P2 and P9 duplicated work-in-progress while the agent continued independently. Coarse-grained interventions (appropriation, opportunistic takeover) validate direction, where P4 and P7 appropriating outputs meant “this works, continue autonomously.” Rather than concurrent interactions as the same signal, agents should interpret each granularity level as calibrated feedback: fine-grained edits demand in-context adaptation, mid-level takeovers signal task redistribution, and coarse-grained appropriation warrants increased autonomy.
@@
DI3. Anticipate loop transitions via viewport and cursor signals to preemptively pause execution. Rather than only reacting after users intervene, agents should track the user’s viewport focus and cursor proximity to predict imminent interventions and pause proactively, minimizing wasted effort and allowing users to redirect execution before unwanted work is completed. When users center their viewport on the agent’s work area while moving their cursor toward its output, this typically signals C or D intervention within seconds (P2, P4, P10); when users maintain a peripheral viewport with the cursor in distant workspace areas, agents can proceed without artificial pauses. When detecting imminent intervention, agents should pause at the current operation boundary and surface the intermediate state, improving execution efficiency (P1 and P6), reducing correction costs (P10), and preserving user agency over the final output (P3, P5, and P9). Also, sustained observation without loop transition signals imminent intervention, prompting agents to prepare an intermediate state rather than proceed autonomously.
@@
DI4. Enable direct manipulation of agent presence as unambiguous collaboration intent signals. While the third implication focuses on agents inferring intent from indirect user information, direct manipulation of agent presence could provide explicit signals. Participants (P1, P7-P9) stated a strong desire for such controls, suggesting they would move the agent away from their work area to signal “work independently elsewhere” H, or position it near specific elements to indicate “focus your attention here” (guiding toward C or D engagement). These spatial manipulations within the shared workspace enable users to clearly communicate when (now or later), where (which work area), and how (proximity suggesting collaboration or distance suggesting independence) they expect the agent to work.
@@
DI5. Drive proactive agent behavior by continuously accumulating and acting on diverse collaboration signals. Current agents operate exclusively in response to user requests, and while recent proactive agents exist, they typically rely on simple heuristics such as detecting prolonged user inactivity (Pu et al., 2025), rather than a rich collaboration context. Our findings suggest that diverse collaboration signals could inform the timing, content, and approach of proactive agent behavior: loop transitions indicating confidence or concern, observation patterns revealing engagement levels, concurrent interventions showing valued directions, and direct canvas manipulations communicating preferred work approaches. By integrating these cues, agents can anticipate user needs beyond current requests, recognizing when sustained observation signals uncertainty warranting proactive explanation, or detecting when users consistently delegate certain task types to proactively suggest similar work (e.g., “I notice you often delegate color exploration, should I generate variations for this new design?”).
5.2. Leveraging Concurrent Interaction as Behavioral Data for Workstyle Learning
@@
Demonstration-based steering and in-situ co-editing of concurrent interaction provide particularly rich behavioral signals, revealing not just what users want, but how they want work done. When P7 applied fine-grained aesthetic refinements, the specific adjustments communicated quality expectations and stylistic preferences more precisely than verbal instruction could convey. Beyond individual actions, loop transitions themselves carry signal value: users moving from observation to concurrent engagement indicate specific aspects requiring attention, while transitions to hands-off validate the agent’s current approach. Our characterized interaction patterns, triggers, and loop dynamics could therefore enable more accurate user simulation (Kim et al., 2026), providing rich behavioral data about not just what users want, but when, how, and why they intervene, supporting more nuanced workstyle learning and intent alignment in future agentic systems. Such learned behavioral patterns and workstyle could further inform the flexible, automatic definition of user-tailored functions grounded in observed workflows (Lam et al., 2025).
5.3. Reification and Domain-Appropriate Representation as Keys to Extending Concurrent Interaction
While our findings emerged from design work, the principles of concurrent interaction could extend to other domains through two key requirements. First, domain-appropriate execution representations: in text-based domains like coding and writing, agents generate content far faster than humans can comprehend, necessitating multi-scale representations where users can zoom between full generated text and hierarchical summaries, with lightweight annotations (“keep”, “revise”, “delete”) enabling concurrent feedback without disrupting generation flow. In web navigation, concurrent interaction is better supported by allowing users to directly manipulate the agent’s intermediate outputs, such as injecting relevant pages into the agent’s path or appropriating search results as directional signals. Second, reification, which means making the agent’s intermediate work products into manipulable objects rather than ephemeral process states, is essential across all domains. Future work should investigate which domain-specific intermediate outputs are most useful for direct manipulation and how agents should interpret such manipulations as collaboration signals.
5.4. Unique Affordances of Human-Agent Collaboration
@@
While we draw on human-human collaboration theories (Gutwin and Greenberg, 2002; Robertson, 2002; Tang, 1991), human-agent collaboration exhibits unique affordances absent in human teams. Users can interrupt without guilt, redirect without negotiation, and appropriate outputs without attribution concerns, behaviors that would carry significant social costs in human collaboration. Participants exploited this asymmetry freely: P1-P3, P6-P8, and P10 terminated execution when correction felt burdensome, while P5 and P9 appropriated agent outputs without acknowledgment, using agent work as raw material for their own designs. Unlike human collaborators, agents accept abrupt termination, interpret overwriting as feedback, and restart without loss of motivation — these unique affordances of human-agent collaboration enable more experimental and iterative workflows that would be unthinkable with human partners.
@@ 5.5. Limitations and Future Work
Our second study was limited to 10 participants over two days, requiring larger-scale longitudinal validation to understand how interaction patterns evolve as users develop mental models of the agent’s capabilities. Additionally, our ReAct-based agent architecture represents one design choice among many; different architectures (e.g., reflection-based, hierarchical planning, continuous execution) may elicit different interaction patterns, and investigating this could reveal which design elements are critical for effective co-creative work. Finally, agent execution speed warrants investigation: slower execution supports monitoring and comprehension, while faster execution helps users reach outcomes earlier, and future work should explore user-controllable speed that adapts based on task stage, user confidence, and intervention frequency.
@@ 6. Conclusion
This work investigates the transition from sequential delegation to dynamic, concurrent human-agent collaboration in co-creative design. Study 1 revealed that process transparency naturally prompted concurrent intervention, motivating Cleo, which interprets and adapts to user actions during execution. Analyzing 214 turn-level interactions, we identified five action patterns, six triggers, and four enabling factors, synthesized into a decision model of six interaction loops. These findings suggest opportunities for future agents to adapt process visibility, interpret concurrent interventions as workstyle signals, and act proactively on accumulated collaboration context, positioning concurrent interaction not as a replacement for delegation, but as what makes delegation work better.
@@ A.3.3. Prompt: Plan Update Module
A.4. Cleo Agent Pipeline in the First Probe
Appendix B Supplementary Results
All results reported in this section are derived from Study 2.
B.1. Interaction Log Processing and Visualization
B.2. User Action Pattern Visualization
Figure 9. Visual representation of user action patterns.
B.3. Turn-Level Action Category Distributions
[…]
(H) 67 31.31%
(O) 30 14.02%
(O) + (H) 22 10.28%
(O) + (C) 11 5.14%
(O) + (D) 5 2.34%
(O) + (T) 3 1.40%
(O) + (C) + (H) 17 7.94%
(O) + (D) + (H) 12 5.61%
(O) + (C) + (D) 12 5.61%
(O) + (H) + (T) 3 1.40%
(O) + (C) + (D) + (H) 19 8.88%
(O) + (D) + (H) + (T) 4 1.87%
(O) + (C) + (D) + (T) 3 1.40%
(O) + (C) + (D) + (H) + (T) 6 2.80%
(O) + (C) + (D) + (H) + (T) 6 2.80%
Table 4. Turn-level distribution of action category combinations across 214 turns.